OpenStudy (anonymous):

If P= 1.2w/h^2 Solve the formula for w Pls help I'm clueless

OpenStudy (anonymous):

P = 1.2w/H^2 Since you want to solve for w, begin to isolate that specified variable: P * H^2 = 1.2 w After this step, divide the left side of the equation by 1.2 and flip the equation around to end up with your final answer: (P*H^2)/1.2 = w w = (P*H^2)/1.2 Hope this helps!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np :)